Why is 66/8 the answer to the following math problem
15/-2 - -3/-4. how do I add the numberators in order to get 66 and then simplify to lowest terms which will be 33/4

Answer by Alan3354(69443)

Why is 66/8 the answer to the following math problem
15/-2 - -3/-4. how do I add the numberators in order to get 66 and then simplify to lowest terms which will be 33/4
----------------
15/-2 - -3/-4
-------
15/2 + 3/4 = 60/8 + 6/8 = 66/8
= 33/4
----------
I'm not sure about all the minus signs you have, tho.
----------
15/-2 - -3/-4
-15/2 - 3/4 = -60/8 - 6/8
= -33/4